if sin a + sin b =c cosa + cosb=d then cos(a+b) is

Grade:11

1 Answers

Nishant Vora IIT Patna
9 years ago
sin a + sin b =c….….(1)
cosa + cosb=d….…..(2)

Cos(A+B) = CosACosB - SinASinB

Square and add equation (1) and (2)
2 + 2 sina sinb + 2 cosa cosb = c2 + d2
2 + 2 cos (a+b) = c2 + d2
From here find out cos(a+b) = ?